# A Comprehensive Survey on Graph Reduction: Sparsification, Coarsening, and Condensation

### v1

@@ Abstract
Many real-world datasets can be naturally represented as graphs, spanning a wide range of domains. However, the increasing complexity and size of graph datasets present significant challenges for analysis and computation. In response, graph reduction techniques have gained prominence for simplifying large graphs while preserving essential properties. In this survey, we aim to provide a comprehensive understanding of graph reduction methods, including graph sparsification, graph coarsening, and graph condensation. Specifically, we establish a unified definition for these methods and introduce a hierarchical taxonomy to categorize the challenges they address. Our survey then systematically reviews the technical details of these methods and emphasizes their practical applications across diverse scenarios. Furthermore, we outline critical research directions to ensure the continued effectiveness of graph reduction techniques, as well as provide a comprehensive paper list at https://github.com/ChandlerBang/awesome-graph-reduction. We hope this survey will bridge literature gaps and propel the advancement of this promising field.
@@ 1 Introduction
Graph-structured data has become ubiquitous in various domains, ranging from social networks and biological systems to recommendation systems and knowledge graphs (Fan et al., 2019; Wu et al., 2022b, 2018; Shi and Weninger, 2017; Wang et al., 2021). The inherent relational structure of graph data makes it a powerful representation for modeling complex interactions and dependencies. Moreover, with the rise of graph machine learning techniques, especially graph neural networks (GNNs) (Kipf and Welling, 2016; Wu et al., 2020), the utilization of graph datasets has seen unprecedented growth, leading to advancements in tasks such as node classification, link prediction, graph classification, and graph generation (Zhou et al., 2020; Ma and Tang, 2021).
@@
Connection to Existing Surveys. In contrast to previous reviews on graph reduction (Liu et al., 2018; Interdonato et al., 2020; Shabani et al., 2023), our study offers a comprehensive overview of the emerging field of graph condensation and presents a unified framework that bridges graph condensation with conventional graph reduction techniques. Additionally, our survey explores synergies between graph reduction and graph neural networks, an aspect rarely covered in existing surveys. Furthermore, our work shares connections with recent surveys on dataset distillation (Geng et al., 2023; Sachdeva and McAuley, 2023), while they predominantly focus on condensation methods applied to image data.
@@
Figure 2: Taxonomy of existing graph reduction methods.
@@ A Unified Framework of Graph Reduction
Given a graph $G=(\mathcal{V},\mathcal{E})$, graph reduction outputs a graph ${G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ which contains $N^{\prime}$ nodes and $E^{\prime}$ edges, subject to $N^{\prime}<N$, or $E^{\prime}<E$ edges. The reduced graph $G^{\prime}$ preserves the desired information of the original graph $G$. This process can be understood as finding a graph $G^{\prime}$ that minimizes a loss function $\mathcal{L}$ explicitly or implicitly, which measures the difference between ${\bf G}$ and ${\bf G}^{\prime}$ in terms of certain information:
@@
Given a graph $G=({\bf A},{\bf X})$, graph coarsening outputs $G^{\prime}=({\bf A}^{\prime},{\bf X}^{\prime})$ containing $N^{\prime}$ super-nodes (clusters) and $E^{\prime}$ super-edges, where $N^{\prime}<N$. It requires finding a surjective mapping from the original graph to a coarse graph, which can be formulated by a one-hot matrix ${\bf C}\in\{0,1\}^{N\times N^{\prime}}$ that assigns nodes to super-nodes. We further define the reverse assignment matrix ${\bf P}=\operatorname{rowNormalize}({\bf C}^{\top})$. Then the coarse graph is usually constructed by ${\bf A}^{\prime}={\bf C}^{\top}{\bf A}{\bf C},{\bf X}^{\prime}={\bf P}{\bf X}$ with the Laplacian matrix being ${\bf L}^{\prime}={\bf C}^{\top}{\bf L}{\bf C}$.
@@
Given a graph $\mathcal{T}=(\mathbf{A},\mathbf{X},\mathbf{Y})$, with ${\bf Y}$ being the node labels, graph condensation aims to learn a small-synthetic graph dataset $\mathcal{S}=(\mathbf{A^{\prime}},\mathbf{X^{\prime}},\mathbf{Y^{\prime}})$, where $N^{\prime}<N$, such that a GNN trained on $\mathcal{S}$ obtains a comparable performance to the one trained on $\mathcal{T}$.
Table 2: General qualitative comparison of graph reduction methods. “Scalability”: the ability to scale up to large graphs, “Interpratability”: the existence of correspondence between nodes in the original and reduced graphs, “Label Utilization“: the reliance on the label information.
Strategy Interpretability Label Utilization Objective Output
Sparsification $\times$ Property Preservation Sampled graph
Coarsening $\times$ Property/Performance Preservation Supergraph
Condensation $\times$ Performance Preservation Synthetic graph
@@ 3.1 Graph Sparsification
Graph sparsification, as an intuitive method for graph reduction, involves the selection of crucial edges or nodes based on specific criteria. Conventional approaches typically focus on preserving specific graph properties, such as spectrum and centrality. With the increasing popularity of GNNs, there is a growing array of methods aimed at maintaining the quality of node representations. Consequently, we categorize existing techniques into two groups based on their preservation goals: those focused on preserving graph properties and those dedicated to maintaining model performance.
@@ 3.1.1 Preserving Graph Properties
Spanners. Althöfer et al. (1993) first develop an algorithm to obtain spanners in graphs. It starts with an empty graph defined on the original node set and adds edges from the original graph only if their weight is smaller than the current distance between their connected nodes in the reduced graph. They also prove that every weighted graph has a $2t$-spanner with $O(N^{1+1/t})$ edges.  Baswana and Sen (2003) tighten this upper bound to $(2t-1)$ with a linear time algorithm that merely explores the edges in the neighborhood of a node. Furthermore, by defining a reinforcement learning process and adjusting reward functions to the preservation of pairwise distance, SparRL (Wickman et al., 2022) outperforms all conventional graph sparsification approaches.
Sparsifiers. One representative sparsifier is called Twice Ramanujan Sparsifier (TRS) (Batson et al., 2009), which prove that for every $\epsilon\in(0,1)$ and every undirected graph $G$, there exists a weighted graph $G^{\prime}$ with at most $(N-1)/2$ edges such that $G^{\prime}$ is the $(1+\epsilon)$-sparsifier of $G$ with high probability. This approach presents an algorithm for deriving $G^{\prime}$ by decomposing the graph into subgraphs with high conductance, calculating pairwise effective resistance (ER) (Spielman and Srivastava, 2008), and sampling edge based on normalized ER as probabilities. Then the edges in the reduced graph are reweighted as the probabilities. Furthermore,  Lee and Sun (2018) present an almost-linear time algorithm for constructing such a sparsifier. Previous studies typically necessitate alterations to edge weights as part of the reweighting process. Anderson et al. (2014) address the sparsifier problem conditioned by keeping the original edge weights, employing a method of unweighted column selection. Since most theories are constructed upon the undirected graph, Chung (2005) first extends them into directed graphs by first symmetrizing the graph Laplacians. Different from the above methods that only cut edges, Feng (2016) first finds an extremely sparse subgraph – low-stretch spanning tree, and recovers small portions of off-tree edges to further approximate the spectrum.
@@ 3.1.2 Preserving Performance
Many methods employ model-free heuristics as the scoring strategy, which calculate the score with metrics derived from the graph structure. For example, Salha et al. (2019) use $k$-core decomposition to find interconnected subgraphs with different density index $k$. By treating subgraphs corresponding to high values of $k$ as the reduced graph, they effectively circumvent the computational demands associated with calculating node embeddings for large graphs. Similarly, Brandeis et al. (2020) utilize this framework for reducing the graph but generate the node embeddings with different methods. Furthermore, recent work by Razin et al. (2023) highlights that the ability to model node interactions is primarily determined by the model-free metrics walk index (WI), which refers to the number of walks originating from the boundary of the partition. Consequently, unimportant edges are removed based on sorted WI values. Although these metrics offer insights from a certain perspective of the graph, they might not be compatible with the downstream models and tasks.
In contrast, recent years have also witnessed many model-based scoring methods, which utilize a parameterized model to calculate the score. For instance, Jin et al. (2022b) adopt coreset methods (Welling, 2009; Sener and Savarese, 2018) to select nodes based on their embeddings from a trained GNN model. Apart from these general scoring methods which can be used for any modality, recent works on the interpretability of GNN, e.g., GNNexplainer (Ying et al., 2019) can also be related to graph sparsification. IGS (Li et al., 2023a) sparsifies a graph based on edge importance obtained from GNNexplainer and feeds the sparsified graph into the next iteration.
Other sparsification methods belong to the graph structure learning (Zheng et al., 2020; Chen et al., 2021; Liu et al., 2023a; Jin et al., 2023) which iteratively optimize the sparsification process by interaction with GNN models. For example, Zheng et al. (2020) learn a sparsified graph structure by neighbor sampling and reparameterization trick (Jang et al., 2016) during GNN training; UGS (Chen et al., 2021) and CGP (Liu et al., 2023a) simultaneously prune the graph adjacency matrix and the GNN weights.
@@ 3.2.1 Reconstruction-Based Methods
Spectral coarsening. Different from spatial methods, spectral coarsening methods compare the ${\bf L}_{l}$ and $\bf L$ by comparing their eigenvalues or eigenvectors. The lifted Laplacian matrix is defined as ${\bf L}_{l}={\bf P}^{\top}{\bf L}^{\prime}{\bf P}$ (Kumar et al., 2023). Loukas (2019); Loukas and Vandergheynst (2018) propose restricted spectral approximation and derive a relaxed evaluation called Relative Eigenvalue Error (REE) defined as $\operatorname{REE}=\sum_{i=1}^{k}|\lambda_{i}-\lambda^{\prime}_{i}|/\lambda_{i}$ , where $\lambda_{i}$ and $\lambda^{\prime}_{i}$ are the top-$k$ eigenvalues of the matrices ${\bf L}$ and ${\bf L}^{\prime}$, respectively. Note that they use ${\bf L}^{\prime}$ instead of ${\bf L}_{l}$ because the comparison of eigenvalues does not require the alignment of the sizes. They also give the theoretical guarantee of greedy edge contraction approaches, where different edge scoring methods can be used including Heavy Edge (Dhillon et al., 2007), Algebraic Distance (Chen and Safro, 2011), Affinity (Livne and Brandt, 2012) and Local Variation (LV) (Loukas, 2019). Some works hold that the edge weights can be further optimized after edge contraction. For example,  Zhao et al. (2018) scale the edge weights by stochastic gradient descent to further align the eigenvalues after coarsening. In addition, some endeavor is made for lossless coarsening, e.g., Navlakha et al. (2008); Khan et al. (2015) keep the correction set recording the missed edges during edge contraction.
@@ 3.2.2 Reconstruction-Free Methods
Despite the proliferation of reconstruction-based methods, other approaches do not rely on the reconstruction while still keeping the key information. Kron reduction (Dorfler and Bullo, 2012) is initially developed to address challenges in electrical networks, specifically to simplify resistance networks while maintaining the pairwise ER. This method calculates the coarse graph Laplacian by

### v4

@@ Abstract
Many real-world datasets can be naturally represented as graphs, spanning a wide range of domains. However, the increasing complexity and size of graph datasets present significant challenges for analysis and computation. In response, graph reduction, or graph summarization, has gained prominence for simplifying large graphs while preserving essential properties. In this survey, we aim to provide a comprehensive understanding of graph reduction methods, including graph sparsification, graph coarsening, and graph condensation. Specifically, we establish a unified definition for these methods and introduce a hierarchical taxonomy to categorize the challenges they address. Our survey then systematically reviews the technical details of these methods and emphasizes their practical applications across diverse scenarios. Furthermore, we outline critical research directions to ensure the continued effectiveness of graph reduction techniques, as well as provide a comprehensive paper list at https://github.com/Emory-Melody/awesome-graph-reduction. We hope this survey will bridge literature gaps and propel the advancement of this promising field.
@@ 1 Introduction
Graph-structured data has become ubiquitous in various domains, ranging from social networks and biological systems to recommendation systems and knowledge graphs (Fan et al., 2019; Wu et al., 2022b, 2018; Shi and Weninger, 2017; Wang et al., 2021). The inherent relational structure of graph data makes it a powerful representation for modeling complex interactions and dependencies. Moreover, with the rise of graph machine learning techniques, especially graph neural networks (GNNs) (Kipf and Welling, 2016; Wu et al., 2020; Ma and Tang, 2021), the utilization of graph datasets has seen unprecedented growth, leading to advancements in fields such as chemistry (Reiser et al., 2022), biology (Gligorijević et al., 2021), and epidemiology (Liu et al., 2024).
@@
Figure 2: Illustration of key differences among three strategies of graph reduction: Graph sparsification selects significant nodes and edges while discarding others, graph coarsening groups and aggregates similar nodes and edges to construct a smaller graph, and graph condensation learns a synthetic graph from scratch.
@@
Connection to Existing Surveys. In contrast to previous reviews on graph reduction (Liu et al., 2018; Interdonato et al., 2020; Shabani et al., 2023; Chen et al., 2022), our study offers a comprehensive overview of the emerging field of graph condensation and presents a unified framework that bridges graph condensation with conventional graph reduction techniques. Compared with the latest survey on graph learning acceleration (Zhang et al., 2023a) which also categorizes graph reduction algorithms into those as mentioned above three, our work uses more general definitions and discusses more related methods on each category in a wider scope. Additionally, our survey explores synergies between graph reduction and GNNs, an aspect rarely covered in existing surveys. Also, some data-centric graph learning surveys (Zha et al., 2023; Zheng et al., 2023a)include discussions on graph reduction but we offer a more detailed and thorough examination of reduction techniques. Furthermore, our work shares connections with recent surveys on dataset distillation (Geng et al., 2023; Sachdeva and McAuley, 2023), while they majorly focus on condensation methods applied to image data.
@@ A Unified Framework of Graph Reduction
Given a graph $G=(\mathcal{V},\mathcal{E})$, graph reduction outputs a graph ${G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ which contains $N^{\prime}$ nodes and $E^{\prime}$ edges, subject to $N^{\prime}<N$, or $E^{\prime}<E$ edges. The reduced graph $G^{\prime}$ preserves the desired information of the original graph $G$. This process can be understood as finding a graph $G^{\prime}$ that minimizes a loss function $\mathcal{L}$ explicitly or implicitly, which measures the difference between $G$ and $G^{\prime}$ in terms of certain information:
@@
Given a graph $G=({\bf A},{\bf X})$, graph coarsening outputs $G^{\prime}=({\bf A}^{\prime},{\bf X}^{\prime})$ containing $N^{\prime}$ super-nodes and $E^{\prime}$ super-edges, where $N^{\prime}<N$. It requires finding a surjective mapping from the original graph to a coarse graph, which can be formulated by a one-hot matrix ${\bf C}\in\{0,1\}^{N\times N^{\prime}}$ that assigns nodes to super-nodes. We further define the reverse assignment matrix ${\bf P}=\operatorname{rowNormalize}({\bf C}^{\top})$. Then the coarse graph is usually constructed by ${\bf A}^{\prime}={\bf C}^{\top}{\bf A}{\bf C},{\bf X}^{\prime}={\bf P}{\bf X}$ with the Laplacian matrix being ${\bf L}^{\prime}={\bf C}^{\top}{\bf L}{\bf C}$.
@@
Given a graph $\mathcal{T}=(\mathbf{A},\mathbf{X},\mathbf{Y})$, with ${\bf Y}$ being the node labels, graph condensation aims to learn a small-synthetic graph dataset $\mathcal{S}=(\mathbf{A^{\prime}},\mathbf{X^{\prime}},\mathbf{Y^{\prime}})$, where $\mathcal{S}$ contains learnable parameters, and $N^{\prime}<N$, such that a GNN trained on $\mathcal{S}$ obtains a comparable performance to the one trained on $\mathcal{T}$.
Table 2: General qualitative comparison of graph reduction methods. “Scalability”: the ability to scale up to large graphs, “Interpretability”: the existence of correspondence between nodes in the original and reduced graphs, “Label Utilization“: the reliance on label information.
Strategy Interpretability Label Utilization Objective Output Remark
Sparsification $\times$ Property/Performance Preservation Subgraph Select subgraphs with maximal information
Coarsening $\times$ Property/Performance Preservation Supergraph Merge close nodes
Condensation $\times$ Performance Preservation Synthetic graph Generates small graphs with high utility
@@ 3.1 Graph Sparsification
Graph sparsification, as an intuitive method for graph reduction, involves selecting essential edges or nodes based on specific criteria and then constructing a smaller graph from these selected elements. Conventional approaches typically focus on preserving specific graph properties, such as spectrum and centrality. With the increasing popularity of GNNs, there is a growing array of methods aimed at maintaining the quality of node representations. Consequently, we categorize existing techniques into two groups based on their preservation goals: those focused on preserving graph properties and those dedicated to maintaining model performance.
@@ 3.1.1 Preserving Graph Properties
Spanners. Althöfer et al. (1993) first develop an algorithm named SPANNER to obtain spanners in graphs. It starts with an empty graph defined on the original node set and adds edges from the original graph only if their weight is smaller than the current distance between their connected nodes in the reduced graph. They also prove that every weighted graph has a $2t$-spanner with $O(N^{1+1/t})$ edges.  Baswana and Sen (2003) tighten this upper bound to $(2t-1)$ with a linear time algorithm that merely explores the edges in the neighborhood of a node. Furthermore, by defining a reinforcement learning process and adjusting reward functions to the preservation of pairwise distance, SparRL (Wickman et al., 2022) outperforms all conventional graph sparsification approaches.
Figure 3: Taxonomy of existing graph reduction methods.
Sparsifiers. One representative sparsifier is called Twice Ramanujan Sparsifier (TRS) (Batson et al., 2009), which prove that for every $\epsilon\in(0,1)$ and every undirected graph $G$, there exists a weighted graph $G^{\prime}$ with at most $(N-1)/2$ edges such that $G^{\prime}$ is the $(1+\epsilon)$-sparsifier of $G$ with high probability. This approach presents an algorithm for deriving $G^{\prime}$ by decomposing the graph into subgraphs with high conductance, calculating pairwise effective resistance (ER) (Spielman and Srivastava, 2008), and sampling edge based on normalized ER as probabilities. Then the edges in the reduced graph are reweighted as the probabilities. Furthermore,  Lee and Sun (2018) presents an almost-linear time algorithm for constructing such a sparsifier. Previous studies typically necessitate alterations to edge weights as part of the reweighting process. Anderson et al. (2014) address the sparsifier problem conditioned by keeping the original edge weights, employing a method of unweighted column selection. Since most theories are constructed upon the undirected graph, Chung (2005) first extends them into directed graphs by first symmetrizing the graph Laplacians. Different from the above methods that only cut edges, Feng (2016) first finds an extremely sparse subgraph – low-stretch spanning tree, and recovers small portions of off-tree edges to further approximate the spectrum. GSGAN (Wu and Chen, 2020) designs a reward function to guide the graph generator in a generative adversarial network creating random walks. These random walks are finally combined to form a smaller edge set that is effective for community detection.
@@ 3.1.2 Preserving Performance
Many methods employ model-free heuristics as the scoring strategy, which calculate the score with metrics derived from the graph structure. For example, Salha et al. (2019) use $k$-core decomposition to find interconnected subgraphs with different density index $k$. By treating subgraphs corresponding to high values of $k$ as the reduced graph, they effectively circumvent the computational demands associated with calculating node embeddings for large graphs. Similarly, CoreWalk (Brandeis et al., 2020) utilizes this framework for reducing the graph but generate the node embeddings with different methods. Furthermore, recent work WIS (Razin et al., 2023) highlights that the ability to model node interactions is primarily determined by the model-free metrics walk index (WI), which refers to the number of walks originating from the boundary of the partition. Consequently, unimportant edges are removed based on sorted WI values. Although these metrics offer insights from a certain perspective of the graph, they might not be compatible with the downstream models and tasks.
In contrast, recent years have also witnessed many model-based scoring methods, which utilize a parameterized model to calculate the score. For instance, Jin et al. (2022b) adopt coreset methods (Welling, 2009; Sener and Savarese, 2018) to select nodes based on their embeddings from a trained GNN model. Apart from these general scoring methods which can be used for any modality, recent works on the interpretability of GNN, e.g., GNNexplainer (Ying et al., 2019) can also be related to graph sparsification. For example, IGS (Li et al., 2023a) sparsifies a graph based on edge importance obtained from GNNexplainer and feeds the sparsified graph into the next iteration. Other methods fall under graph structure learning (Zhu et al., 2021), which can be viewed as renders the scoring of edges learnable. For example, Zheng et al. (2020) learn a sparsified graph structure by neighbor sampling according to the and reparameterization trick (Jang et al., 2016) during GNN training; UGS (Chen et al., 2021) and CGP (Liu et al., 2023a) simultaneously prune the elements in graph adjacency matrix (edges) and the GNN weights to reduce the graph and lighten the model. Recognizing that UGS fails to preserve topology, GST (Zhang et al., 2024a) is proposed to combine semantic and topological information during sparsification. Differing from the iterative approaches described above, TEDDY (Seo et al., 2024) introduces a one-shot reduction strategy. This strategy begins by sparsifying the graph using only degree information, followed by training a lightweight model through knowledge distillation and Projected Gradient Descent (PGD) on the $l_{0}$ ball. However, these structure learning methods are closely tied to specific selection models or sparsified GNN models. In other words, these graphs are merely byproducts of graph structure learning, even though they have the potential to serve as reduced inputs for other algorithms. The transferability and universality of their sparsified graphs to other graph algorithms remain inadequately explored.
There are a few works that do not fall into the scoring-and-selection strategy and instead adopt a holistic approach to selection. For example, OTC (Garg and Jaakkola, 2019) achieves selection by minimizing the optimal transport distance between an initial node distribution supported on $\mathcal{V}$ and a target distribution supported on a subset $\mathcal{V}^{\prime}$ selected via the associated transport problem (Peyré et al., 2019). The reduced graph is then obtained by restricting the original graph to nodes in $\mathcal{V}^{\prime}$ and their incident edges.
@@ 3.2.1 Reconstruction-Based Methods
Spectral coarsening. Different from spatial methods, spectral coarsening methods compare the ${\bf L}_{l}$ and $\bf L$ by comparing their eigenvalues or eigenvectors. The lifted Laplacian matrix is defined as ${\bf L}_{l}={\bf P}^{\top}{\bf L}^{\prime}{\bf P}$ (Kumar et al., 2023). Loukas (2019); Loukas and Vandergheynst (2018) propose restricted spectral approximation and derive a relaxed evaluation called Relative Eigenvalue Error (REE) defined as $\operatorname{REE}=\sum_{i=1}^{k}|\lambda_{i}-\lambda^{\prime}_{i}|/\lambda_{i}$ , where $\lambda_{i}$ and $\lambda^{\prime}_{i}$ are the top-$k$ eigenvalues of the matrices ${\bf L}$ and ${\bf L}^{\prime}$, respectively. Note that they use ${\bf L}^{\prime}$ instead of ${\bf L}_{l}$ because the comparison of eigenvalues does not require the alignment of the sizes. They also give the theoretical guarantee of greedy pairwise contraction approaches, where different node pair scoring methods can be used including Heavy Edge (Dhillon et al., 2007), Algebraic Distance (Chen and Safro, 2011), Affinity (Livne and Brandt, 2012) and Local Variation (LV) (Loukas, 2019). Some works hold that the edge weights can be further optimized after pairwise contraction. For example,  Zhao et al. (2018) scale the edge weights by stochastic gradient descent to further align the eigenvalues after coarsening. In addition, some endeavor is made for lossless coarsening, e.g., Navlakha et al. (2008); Khan et al. (2015) keep the correction set recording the missed edges during pairwise contraction.
@@ 3.2.2 Reconstruction-Free Methods
Despite the proliferation of reconstruction-based methods, other approaches do not rely on the reconstruction while still keeping the key information.  (Itzkovitz et al., 2005) coarsen complex engineered and biological networks into smaller and more comprehensible versions, where nodes represents probabilistically generalized network motifs in the original network. To analyze social networks with diverse attributes and relations, SNAP (Tian et al., 2008) produces a summary graph where every node inside a super node has the same values for selected attributes, and is adjacent to nodes with the same selected relations. k-snap (Tian et al., 2008) relaxes this homogeneity requirement and allows users to control the resolutions of summaries. AGSUMMARY (Wu et al., 2014) utilizes the Minimum Description Length principle to design a cost function and compute an optimal summary by neighborhood greedy strategy. Since the former two methods only apply to discrete attributes, CANCEL (Zhang et al., 2010) relaxes this condition to continual ones with adaptive cutoffs and proposes a comprehensive metric named interestingness. To extend the above methods focusing only on one task, CoarseNet (Purohit et al., 2014) tries to find a succinct representation of the original network that preserves important diffusive characters, which can be applied to both influence maximization and propagation patterns detection tasks. To flexibly achieve higher performance among different tasks, Netgist (Amiri et al., 2018) defines a task-based graph summarization problem and uses RL to create a flexible framework for learnable node merging policies.
3.2.3 Remarks on Graph Coarsening
Graph Coarsening in GNNs. There are growing numbers of works that combine coarsening with GNNs. To mimic the pooling layer in the convolutional neural network, Such et al. (2017) make the mapping matrix learnable and produce a pooled graph reduced to fewer nodes layer by layer. $k$-MIS (Bacciu et al., 2023) introduces a unifying interpretation of pooling in both regular and graph data with a controllable equispaced graph coarsening mechanism. Borrowing from graph covers and k-plexes (pseudo-cliques where each node can miss up to $k$ links), KPlexPool Bacciu et al. (2021) is developed as a non-parametric pooling technique that can generalize effectively to graphs with varying natures and connectivity patterns.
While coarsening methods for pooling are mostly tailored for graph classification, recent years have also seen the development of coarsening methods for node classification. For instance, SCAL (Huang et al., 2021) first trains a GNN model in a graph coarsened by LV, with super node label defined as ${\bf Y}^{\prime}=\operatorname{argmax}({\bf P}{\bf Y})$ and then directly uses this model to inference. Following the same framework, Generale et al. (2022) define the relation summarization and uses R-GCN (Schlichtkrull et al., 2018) for knowledge graphs. CONVMATCH (Dickens et al., 2023) merges nodes that are equivalent or similar w.r.t. the GCN convolution operation. Similarly,  Buffelli et al. (2022) match the node embeddings output by GNNs among graphs in different coarsening ratios to deal with one of data shift problem — size shift.
Some works employ coarsening techniques to improve the scalability when training GNNs in large-scale graphs. VNG (Si et al., 2023) highlights the ongoing challenge of efficiently deploying a GNN model in online applications when edges exist between testing nodes and training nodes. To tackle this issue, they match the forward propagation to obtain the mapping matrix ${\bf C}$ by applying weighted k-means. Recognizing that only a subset of nodes (target nodes) requires analysis in large-scale web graphs, Graph Skeleton (Cao et al., 2024) focuses reducing the background nodes while preserving the classification performance of target nodes. It merges background nodes sharing the same distance to target nodes, as well as target nodes and their affiliated nodes—that is, background nodes linked exclusively to a single target node.
Some methods do not change the node features but the paths between each pair of target nodes. For example, Kron reduction (Dorfler and Bullo, 2012) is initially developed to address challenges in electrical networks, specifically to simplify resistance networks while maintaining the pairwise ER. This method calculates the coarse graph Laplacian ${\bf L^{\prime}}$ by Schur complement